OpenStudy (anonymous):

Write the equation of a parabola with a vertex at (0, 0) and a focus at (0, -5).

OpenStudy (anonymous):

A. x=-1/20y^2 B. y=-1/20x^2 C. y=1/20x^2 D. x=1/20y^2

OpenStudy (tkhunny):

The all have Vertex at (0,0). That's no help.

OpenStudy (tkhunny):

Focus is (0,-5). This means, in relation to the vertex, that it is a downward opening parabola. Throw out everything with y^2 and everything with a positive coefficient. That about does it.

OpenStudy (anonymous):

thanks!!!!

OpenStudy (tkhunny):

Do another one. Let's see if you're getting it.